OpenStudy (anonymous):

Simplify. x8 ÷ x4 · x2

OpenStudy (anonymous):

if you mean (x8/x4) * x2 then the answer is 4x (as long as x does not equal 0)

OpenStudy (anonymous):

x^8 divided by x^4 multiplied by x^2

OpenStudy (anonymous):

oh that changes the question a lot: \[x ^{8} \div x^4 * x^2 = x^6\]

jhonyy9 (jhonyy9):

x*8 x*8 ---------------- = ------- = x*2 x*(4) multiplied x*(2) x*6